Two similar sonometer wires of same material under the same tension produce two beats per second.The length of one wire is 50 cm and that of the other is 50.1cm.calculate the frequencies of the two wires.

Grade:12th pass

1 Answers

Diana
4 years ago
n1-n2=2.......(1)
l1 = 50 cm
l2 = 50.1 cm
 
As the tensions and the material of wire are the same 
     n1 l1 = n2l2 
     Or n1/n2= l2/l1
    n1=( 50.1/50 ) n2 ......(2)
 
 
 Sub. in (1) 
(50.1/50) n2 - n2 = 2
n2 = 1000 Hz
N1= 1000 + 2
 n1= 10002 hz